OpenStudy (anonymous):

solve the system using the substitution method. 2x-y=1 8x+4y=6 i got (0.625, 0.25) am i right

OpenStudy (anonymous):

@pleasehelpineedhelp Would you like to show your work so somebody can figure out your mistake?

OpenStudy (anonymous):

Um firstly, get one common variable at the top, so you would for example substract that 2x from 2x-y=1 and than you would have -y=-2x, divide that negative so you get y=2x. Substitute that y with 4y from 8x+4y=6, so it becomes 8x+4(2x)=6. Your answer was incorrect, I hope you can go from there or do you need further steps?